OpenStudy (anonymous):

(1)/(5)=()/(15) find the number with denominator

OpenStudy (anonymous):

do you mean (1/5)=(x/15)? in this case x=3

OpenStudy (anonymous):

(1)/(5)=()/(15) find the number with denominator indicated multiply by 1